OpenStudy (anonymous):

Which of the binomials below is a factor of this trinomial? x2 + 14x + 40 A. x + 10 B. x - 9 C. x2 + 4 D. x + 14

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Factor to (x+4) and (x+10) and (x+10) is the only choice

OpenStudy (maheshmeghwal9):

(x+10)(x+4)= ur question. So what should u choose?

OpenStudy (anonymous):

x2 + 14x + 40=(x+4)(x+10) so answer A

OpenStudy (anonymous):

thanks